OpenStudy (amtran_bus):

If xy^2 = 12 and xy = 4, what is x?

OpenStudy (amtran_bus):

\[x(\frac{ 4 }{ x })^2=12\]

OpenStudy (amtran_bus):

\[x(\frac{ 16 }{ x^2 })=12\]

imqwerty (imqwerty):

you're doing good so far any ideas about the next step? :)

OpenStudy (amtran_bus):

Well I would multiply x and x^2

imqwerty (imqwerty):

yeah

imqwerty (imqwerty):

well it must be multiply x and 1/x^2

OpenStudy (agent0smith):

\[\frac{ x }{ 1 }(\frac{ 16 }{ x^2 })=12\]

imqwerty (imqwerty):

\(\large\cancel {x}\frac{16}{x^{\cancel 2 \color{red}1}}=12\)

OpenStudy (amtran_bus):

ohhh!!!!!!!

OpenStudy (mathstudent55):